Draw all possible random samples of size 3 with replacement from the 

population 15, 24 ,38 ,10 . all sample mean equal to population mean


1
Expert's answer
2022-01-04T16:47:59-0500

Population mean "\mu = \frac{15+24+38+10}{4}= 21.75"

N=4

n=3

Number of samples "= N^n = 4^3 = 64"